### v1

# references.bib

Abstract
In this paper, we consider the resonance problem for the cubic nonlinear Helmholtz equation in the subwavelength regime. We derive a discrete model for approximating the subwavelength resonances of finite systems of high-contrast resonators with Kerr-type nonlinearities. Our discrete formulation is valid in both weak and strong nonlinear regimes. Compared to the linear formulation, it characterizes the soliton-like extra eigenmodes that have recently been experimentally observed.
Keywords. Nonlinear subwavelength physics, nonlinear subwavelength resonance, capacitance matrix formalism, cubic nonlinear Helmholtz equation, soliton-like eigenmode.
AMS Subject classifications. 35P30,35C20,74J20.
1 Introduction and problem formulation
In nanophotonics and nanophononics, the ideal artificial material shall, at subwavelength scales, confine and guide waves, and be able to modify the phase and enhance the amplitude of incoming waves. For doing so, a class of nonlinear artificial materials consisting of nonlinear subwavelength high-contrast resonator building blocks has been introduced. This class of materials has revived the exciting prospect of realizing the long-standing goal of robust transport of waves at subwavelength scales and their amplification in only one direction, opening avenues to channel energy and information transport in quantum computing; see, e.g., [soliton3, rev-acoustics, soliton1, active, soliton2, nonreciprocal].
Although very significant experimental and numerical modeling advances have been achieved in the field of nonlinear subwavelength physics during recent decades, very little is known from a mathematical point of view. As far as we know, properties of nonlinear subwavelength resonances have not yet been analyzed in the mathematical literature. In contrast to the linear response regime studied in depth in [cbms, review], there is a clear lack of deep understanding of the theory of nonlinear subwavelength resonances.
In this work, we lie the basis for nonlinear subwavelength physics. The focus is to develop a theory of nonlinear subwavelength resonances. This is accomplished by using ideas from [feppon2024subwavelength] to the setting of nonlinear systems. It is worth emphasizing that the nonlinearity renders the nonlinear problem much more difficult than the linear one. The total number of subwavelength resonances may be larger than the total number of subwavelength resonators and depends on the strength of the nonlinearities as opposed to the linear systems where the total number of subwavelength resonances is exactly given by the number of resonators. Some soliton-like extra-eigenmodes may exist due to strong enough nonlinearities. These eigenmodes are of profound significance to nonlinear subwavelength physics. They have been experimentally observed in systems of high-contrast resonators where the nonlinearity is of Kerr type [soliton1].
Our main contribution in this work is to provide for the first time a discrete formulation for approximating the subwavelength resonances of finite systems of high-contrast resonators with Kerr-type nonlinearities, which is valid in both weak and strong nonlinear regimes. Compared to the linear formulation, our formulation here characterizes the soliton-like extra eigenmodes. Our main tool for building such a discrete framework is a generalization of the Dirichlet-to-Neumann approach for subwavelength resonances from [feppon2024subwavelength] to the nonlinear setting combined with asymptotic analysis of solutions to nonlinear resonance problems for the Helmholtz equation (with respect to the contrast in the material parameters of the resonators).
Concretely, let $\operatorname{{\mathcal{D}}}\subset\operatorname{{\mathbb{R}^{3}}}3$ be a smooth, bounded domain with connected components $B_{1},\ldots,B_{N}$ that have connected boundaries. Let $\kappa_{0}\in\mathbb{R}$, $\rho_{0}\in\mathbb{R}$ and let $\kappa_{r}:\operatorname{{\mathcal{D}}}\rightarrow\mathbb{R}$, $\rho_{r}:\operatorname{{\mathcal{D}}}\rightarrow\mathbb{R}$ be piecewise constant functions. We consider the following system of equations:
$\left\{\begin{array}[]{rll}\displaystyle\frac{1}{\kappa_{0}}\frac{\partial^{2}u}{\partial t^{2}}-\frac{1}{\rho_{0}}\Delta u&=0&\text{in }\mathbb{R}\times\operatorname{{\mathbb{R}^{3}}}3\setminus\overline{\operatorname{{\mathcal{D}}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\frac{1}{\kappa_{r}(x)}\frac{\partial^{2}u}{\partial t^{2}}-\frac{1}{\rho_{r}(x)}\Delta u&=g(u)&\text{in }\mathbb{R}\times\operatorname{{\mathcal{D}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\frac{1}{\rho_{0}}\left.\frac{\partial u}{\partial\nu}\right|_{+}&\displaystyle=\frac{1}{\rho_{r}(x)}\left.\frac{\partial u}{\partial\nu}\right|_{-}&\text{on }\mathbb{R}\times\partial\operatorname{{\mathcal{D}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\left.u\right|_{+}&=\left.u\right|_{-}&\text{on }\mathbb{R}\times\partial\operatorname{{\mathcal{D}}},\end{array}\right.$ (1)
where $u$ is subject to an outgoing radiation condition, $\nu$ is the outward normal to $\operatorname{{\mathcal{D}}}$, and $g$ is a nonlinear operator. In this article, we consider mainly
$g(u)=b\left|\frac{\partial u}{\partial t}\right|^{2}\frac{\partial u}{\partial t}\quad\text{and}\quad g(u)=b\left(\frac{\partial u}{\partial t}\right)^{2}\overline{u},$
with $b\in\mathbb{C}$. For given operator $g$, we say that $\omega\in\mathbb{C}$ is a resonance of the system (1), if there exists a non-zero $u:\operatorname{{\mathbb{R}^{3}}}3\rightarrow\mathbb{C}$ such that
$(t,x)\in\mathbb{R}\times\operatorname{{\mathbb{R}^{3}}}3\quad\longmapsto\quad\exp(i\omega t)u(x)\in\mathbb{C}$
is a solution to (1). In the linear case, when $g(u)=0$, resonances are precisely those $\omega\in\mathbb{C}$ such that the scattering problem
$\left\{\begin{array}[]{rll}\displaystyle\frac{1}{\rho_{0}}\Delta u+\frac{\omega^{2}}{\kappa_{0}}u&=0&\text{in }\operatorname{{\mathbb{R}^{3}}}3\setminus\overline{\operatorname{{\mathcal{D}}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\frac{1}{\rho_{r}(x)}\Delta u+\frac{\omega^{2}}{\kappa_{r}(x)}u&=0&\text{in }\operatorname{{\mathcal{D}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\frac{1}{\rho_{0}}\left.\frac{\partial u}{\partial\nu}\right|_{+}&\displaystyle=\frac{1}{\rho_{r}(x)}\left.\frac{\partial u}{\partial\nu}\right|_{-}&\text{on }\partial\operatorname{{\mathcal{D}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\left.u\right|_{+}&=\left.u\right|_{-}&\text{on }\partial\operatorname{{\mathcal{D}}},\end{array}\right.$ (2)
where $u-u_{in}$ is subject to an outgoing radiation condition with the incident wave $u_{in}$ satisfying $\frac{1}{\rho_{0}}\Delta u_{in}+\frac{\omega^{2}}{\kappa_{0}}u_{in}=0$ in $\operatorname{{\mathbb{R}^{3}}}3$, does not have a unique solution or, equivalently, is degenerate. We rewrite the material parameters of (1) and (2) as follows:
${c_{0}}:=\sqrt{\frac{\kappa_{0}}{\rho_{0}}},\quad{c_{r}}:=\sqrt{\frac{\kappa_{r}}{\rho_{r}}},\quad\delta(x):=\frac{\rho_{r}(x)}{\rho_{0}}.$
For simplicity of notation, we assume that $\delta(x)\equiv\delta$ and $c_{r}(x)\equiv c_{r}$ are constant across the different connected components of $\operatorname{{\mathcal{D}}}$. We consider the regime when $\delta\rightarrow 0$ and study those resonances which satisfy $\omega(\delta)\rightarrow 0$ as $\delta\rightarrow 0$. To this end, we assume that ${c_{0}}$ and ${c_{r}}$ are constant as $\delta\rightarrow 0$.
Definition 1.1 (Subwavelength resonance).
Let $\omega:(0,\epsilon)\rightarrow\mathbb{C}$ be a continuous function for some $\epsilon>0$. If $\omega(\delta)$ is a resonance of the system (2) for every $\delta\in(0,\epsilon)$ and if $\omega(\delta)\rightarrow 0$ as $\delta\rightarrow 0$, then we say that $\omega(\delta)$ is a (linear) subwavelength resonance.
Linear subwavelength resonances have been thoroughly studied since the original paper [ihp]; see, for instance, [cbms, review]. It has been found that the leading-order behavior of subwavelength resonances $\omega(\delta)$ as $\delta\rightarrow 0$ are characterized by the so-called generalized capacitance matrix $\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})\in\mathbb{R}^{N\times N}$. In fact, one has the following theorem.
Theorem 1.2 (Fundamental theorem of subwavelength physics [review]).
Let $\operatorname{{\mathcal{D}}}\subset\mathbb{R}^{3}$ be a smooth, bounded domain with well separated connected components $B_{1},\ldots,B_{N}$ which are supposed to have connected boundaries. Then there exists a matrix $\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})\in\mathbb{R}^{N\times N}$ depending on $\operatorname{{\mathcal{D}}}$ and $c_{r}$ and there exists an $\epsilon>0$, such that the following correspondence holds
$\begin{array}[]{c}\left\{\begin{array}[]{cc}\omega(\delta):(0,\epsilon)\longrightarrow\mathbb{C}~{}|~{}\omega\text{ is a subwavelength resonance}\end{array}\right\}\vspace{0.2cm}\\ \tilde{\longleftrightarrow}\vspace{0.2cm}\\ \left\{\begin{array}[]{cc}\lambda\in\mathbb{R}~{}|~{}\lambda\text{ is an eigenvalue of }\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})\end{array}\right\},\end{array}$
where $\omega(\delta)$ is associated to the eigenvalue $\lambda$ of $\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})$ such that
$\lvert\omega(\delta)-\sqrt{\lambda}\sqrt{\delta}\rvert=O(\delta)\quad\text{ as}\quad\delta\longrightarrow 0.$
This theorem has been proved using Gohberg-Sigal theory and layer potential techniques. It lays the basis of many results in subwavelength physics. It has been reproduced using the different approach of Dirichlet-to-Neumann operator by Feppon and Ammari [feppon2024subwavelength] for connected $\operatorname{{\mathcal{D}}}=B_{1}$. In this article, we will enlarge the approach of Dirichlet-to-Neumann operators to disconnected $\operatorname{{\mathcal{D}}}=B_{1}\cup\ldots\cup B_{N}$, reproducing Theorem 1.2, in Section 3. Using similar techniques, we will then develop a theory of nonlinear subwavelength resonances in Section 4.
Let $H^{1}(\operatorname{{\mathcal{D}}})$ denote the usual Sobolev space of square-integrable functions whose weak derivative is square integrable in $\operatorname{{\mathcal{D}}}$. We also denote by $H^{\frac{1}{2}}(\partial\operatorname{{\mathcal{D}}})$ the set of traces on $\partial\operatorname{{\mathcal{D}}}$ of functions in $H^{1}(\operatorname{{\mathcal{D}}})$ and let $H^{-\frac{1}{2}}(\partial\operatorname{{\mathcal{D}}})$ be the dual of $H^{\frac{1}{2}}(\partial\operatorname{{\mathcal{D}}})$.
Concretely, the approach is as follows. Let $\mathcal{T}^{k}_{\operatorname{{\mathcal{D}}}}:H^{\frac{1}{2}}(\partial\operatorname{{\mathcal{D}}})\rightarrow H^{-\frac{1}{2}}(\partial\operatorname{{\mathcal{D}}})$ be the Dirichlet-to-Neumann operator associated to the exterior Helmholtz problem
$(\Delta+k^{2})u=0\quad\text{on }\mathbb{R}^{3}\setminus\overline{\operatorname{{\mathcal{D}}}},$
subject to an outgoing radiation condition. This allows to rewrite the system (2) in terms of $\mathcal{T}_{\operatorname{{\mathcal{D}}}}^{\frac{\omega}{{c_{0}}}}$ and $u|_{\operatorname{{\mathcal{D}}}}$ as follows:
$\left\{\begin{array}[]{rll}\displaystyle\Delta u+\frac{\omega^{2}}{{c_{r}}^{2}}u&=0&\text{in }\operatorname{{\mathcal{D}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\frac{\partial u}{\partial\nu}&=\displaystyle\delta\left(\mathcal{T}_{\operatorname{{\mathcal{D}}}}^{\frac{\omega}{{c_{0}}}}[u-u_{in}]+\frac{\partial u_{in}}{\partial\nu}\right)&\text{on }\partial\operatorname{{\mathcal{D}}}.\end{array}\right.$
Formulating this system variationally leads to the following system in terms of $u|_{\operatorname{{\mathcal{D}}}}$ and $\omega$:
$\begin{split}\int_{\operatorname{{\mathcal{D}}}}\nabla u\cdot\nabla\overline{v}\operatorname{dvol}-\frac{\omega^{2}}{{c_{r}}^{2}}\int_{\operatorname{{\mathcal{D}}}}u\overline{v}\operatorname{dvol}-\delta\int_{\partial\operatorname{{\mathcal{D}}}}\mathcal{T}_{\operatorname{{\mathcal{D}}}}^{\frac{\omega}{{c_{0}}}}[u]\overline{v}\operatorname{d\sigma}=\delta\int_{\partial\operatorname{{\mathcal{D}}}}\left(\frac{\partial u_{in}}{\partial\nu}-\mathcal{T}^{\frac{\omega}{v_{0}}}[u_{in}]\right)\operatorname{d\sigma},\\ \quad\forall v\in H^{1}(\operatorname{{\mathcal{D}}}).\end{split}$
Then $\omega$ is a resonance whenever this system is degenerate and does not have a unique solution $u|_{\operatorname{{\mathcal{D}}}}$. Or equivalently $\omega$ is a resonance whenever this system admits a non-zero solution $u|_{\operatorname{{\mathcal{D}}}}$ for $u_{in}\equiv 0$. It was observed in [feppon2024subwavelength] that one can facilitate the analysis and transform this indirect problem of finding characteristic points $\omega$ of the above system into a direct problem, in the case of a connected domain $\operatorname{{\mathcal{D}}}$. By adding $\int_{\operatorname{{\mathcal{D}}}}u\operatorname{dvol}\int_{\operatorname{{\mathcal{D}}}}\overline{v}\operatorname{dvol}$ to the bilinear form on the left and $\int_{\operatorname{{\mathcal{D}}}}\overline{v}\operatorname{dvol}$ to the linear functional on the right one obtains
$\begin{split}\int_{\operatorname{{\mathcal{D}}}}\nabla u\cdot\nabla\overline{v}\operatorname{dvol}+\int_{\operatorname{{\mathcal{D}}}}u\operatorname{dvol}\int_{\operatorname{{\mathcal{D}}}}\overline{v}\operatorname{dvol}-\frac{\omega^{2}}{{c_{r}}^{2}}\int_{\operatorname{{\mathcal{D}}}}u\overline{v}\operatorname{dvol}-\delta\int_{\partial\operatorname{{\mathcal{D}}}}\mathcal{T}_{\operatorname{{\mathcal{D}}}}^{\frac{\omega}{{c_{0}}}}[u]\overline{v}\operatorname{d\sigma}=\int_{\operatorname{{\mathcal{D}}}}\overline{v}\operatorname{dvol},\\ \quad\forall v\in H^{1}(\operatorname{{\mathcal{D}}})\end{split}$ (3)
for $u_{in}\equiv 0$. In the case of sufficiently small $\delta\in\mathbb{R}_{>0}$ and sufficiently small $\omega\in\mathbb{C}$, the bilinear form on the left remains coercive as a small perturbation of the continuous coercive bilinear form [feppon2024subwavelength]
$\begin{split}a_{\omega,\delta}:~{}(u,v)\in H^{1}(\operatorname{{\mathcal{D}}})\times H^{1}(\operatorname{{\mathcal{D}}})\longmapsto\int_{\operatorname{{\mathcal{D}}}}\nabla u\cdot\nabla\overline{v}\operatorname{dvol}+\int_{\operatorname{{\mathcal{D}}}}u\operatorname{dvol}\int_{\operatorname{{\mathcal{D}}}}\overline{v}\operatorname{dvol}.\end{split}$
Then the resonances of (2) are characterized as follows [feppon2024subwavelength]: $\omega$ is a resonance if and only if the unique solution $u$ to the variational problem (3) satisfies
$\int_{\operatorname{{\mathcal{D}}}}u\operatorname{dvol}=1.$ (4)
In the subwavelength regime, $\delta\rightarrow 0$, it is possible to solve the variational problem (3) with the Ansatz
$u_{\omega,\delta}=\sum_{p\geq 0,k\geq 0}\omega^{p}\delta^{k}u_{p,k}.$
Computing $\int_{\operatorname{{\mathcal{D}}}}u_{\omega,\delta}\operatorname{dvol}$ and equating
$\sum_{p\geq 0,k\geq 0}\omega^{p}\delta^{k}\int_{\operatorname{{\mathcal{D}}}}u_{p,k}\operatorname{dvol}=1$
allows to solve for $\omega(\delta)$ and determining the subwavelength resonances in the setting of a connected domain $\operatorname{{\mathcal{D}}}$, see [feppon2024subwavelength, Equation (3.21)].
In this work we generalize the Dirichlet-to-Neumann operator approach to disconnected resonator domains $\operatorname{{\mathcal{D}}}=B_{1}\cup\ldots\cup B_{N}$ (in Section 3) and to nonlinear systems (in Section 4). The main difference is the following. Instead of adding $\int_{\operatorname{{\mathcal{D}}}}u\operatorname{dvol}\int_{\operatorname{{\mathcal{D}}}}\overline{v}\operatorname{dvol}$ to obtain a coercive bilinear form similar to (3), one has to account for the disconnectedness of $\operatorname{{\mathcal{D}}}=B_{1}\cup\ldots\cup B_{N}$ and add $\sum_{i=1}^{N}\int_{B_{i}}u\operatorname{dvol}\int_{B_{i}}\overline{v}\operatorname{dvol}$ instead. In that case one obtains instead of (4) a system of equations accounting for the integral of $u$ on the different connected components $B_{i}$ of $\operatorname{{\mathcal{D}}}$, the details can be found in Subsection 3.1. Similarly to the connected case, one can make the Ansatz $u_{\omega,\delta}=\sum_{p\geq 0,k\geq 0}\omega^{p}\delta^{k}u_{p,k}$ and solve the system of equations asymptotically for subwavelength resonances $\omega(\delta)$. This is presented in Subsection 3.2.
Suppose that $g(u)=\beta\lvert\frac{\partial u}{\partial t}\rvert^{2}\frac{\partial u}{\partial t}$. Then we study the following system of equations:
$\left\{\begin{array}[]{rll}\displaystyle\frac{1}{\kappa_{0}}\frac{\partial^{2}u}{\partial t^{2}}-\frac{1}{\rho_{0}}\Delta u&=0&\text{in }\mathbb{R}\times\operatorname{{\mathbb{R}^{3}}}3\setminus\overline{\operatorname{{\mathcal{D}}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\frac{1}{\kappa_{r}(x)}\frac{\partial^{2}u}{\partial t^{2}}-\frac{1}{\rho_{r}(x)}\Delta u&\displaystyle=\beta\left|\frac{\partial u}{\partial t}\right|^{2}\frac{\partial u}{\partial t}&\text{in }\mathbb{R}\times\operatorname{{\mathcal{D}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\frac{1}{\rho_{0}}\left.\frac{\partial u}{\partial\nu}\right|_{+}&\displaystyle=\frac{1}{\rho_{r}(x)}\left.\frac{\partial u}{\partial\nu}\right|_{-}&\text{on }\mathbb{R}\times\partial\operatorname{{\mathcal{D}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\left.u\right|_{+}&\displaystyle=\left.u\right|_{-}&\text{on }\mathbb{R}\times\partial\operatorname{{\mathcal{D}}},\end{array}\right.$ (5)
subject to an outgoing radiation condition, where we consider the Kerr-like nonlinearity $\beta\lvert\frac{\partial u}{\partial t}\rvert^{2}\frac{\partial u}{\partial t}$ inside the resonator domain $\operatorname{{\mathcal{D}}}$. As in the linear case, we will say that $\omega\in\mathbb{C}$ is a resonance of the system (5) whenever it admits a solution of the form $\exp(i\omega t)u(x)$, where $u:\operatorname{{\mathbb{R}^{3}}}3\rightarrow\mathbb{C}$ is non-zero. This allows to consider the following nonlinear Helmholtz resonance problem:
$\left\{\begin{array}[]{rll}\displaystyle\frac{1}{\rho_{0}}\Delta u+\frac{\omega^{2}}{\kappa_{0}}u&\displaystyle=0&\displaystyle\text{in }\operatorname{{\mathbb{R}^{3}}}3\setminus\overline{\operatorname{{\mathcal{D}}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\frac{1}{\rho_{r}(x)}\Delta u+\frac{\omega^{2}}{\kappa_{r}(x)}u&\displaystyle=i\beta\lvert\omega\rvert^{2}\omega\lvert u\rvert^{2}u&\displaystyle\text{in }\operatorname{{\mathcal{D}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\frac{1}{\rho_{0}}\left.\frac{\partial u}{\partial\nu}\right|_{+}&\displaystyle=\frac{1}{\rho_{r}(x)}\left.\frac{\partial u}{\partial\nu}\right|_{-}&\displaystyle\text{on }\partial\operatorname{{\mathcal{D}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle u|_{+}&\displaystyle=\left.u\right|_{-}&\text{on }\partial\operatorname{{\mathcal{D}}},\end{array}\right.$ (6)
subject to an outgoing radiation condition. As for (2), we will say that $\omega(\delta)$ is a subwavelength resonance, whenever it is a continuous map $\omega:(0,\epsilon)\rightarrow\mathbb{C}$ for some $\epsilon>0$ such that $\omega(\delta)$ is a resonance of (6) for all $\delta\in(0,\epsilon)$ and such that $\omega(\delta)\rightarrow 0$ as $\delta\rightarrow 0$.
The approach to subwavelength resonances in the nonlinear setting is identical to the approach in the linear setting in Section 3, with the exception that one has to consider a nonlinear form $(u,v)\mapsto\tilde{a}_{\omega,\delta}(u,v)$ for the variational formulation of (6), see Subsection 4.1 and that one has to keep track of the amplitudes of the solutions to the variational problem, see Theorem 4.3. The subwavelength resonances are then determined in the same way as for the linear setting, by expanding in $\omega$ and $\delta$ and determining $\omega(\delta)$ such that the characterization of a subwavelength frequency is satisfied, see Subsection 4.2.
This article is organized as follows. In Section 2 we set the notation and the technical prerequisites which will be of use in the later sections. Section 3 is dedicated to the characterization of subwavelength resonances of linear Helmhotz systems with disconnected resonator domain $\operatorname{{\mathcal{D}}}$. It is divided into Subsection 3.1 where the partial differential equation (PDE) system is formulated variationally and the resonances are variationally characterized and into Subsection 3.2 where the asymptotics of subwavelength resonances are computed. Section 4 then builds on top of these techniques and follows a similar structure. First in Subsection 4.1 the system of nonlinear PDE’s is formulated variationally and resonances are characterized. Then in Subsection 4.2 the asymptotics of nonlinear subwavelength resonances are derived. In Section 5 simulations of the leading-order asymptotics of the subwavelength resonances of a nonlinear system of dimers $\operatorname{{\mathcal{D}}}=B_{1}\cup B_{2}$ are presented, demonstrating the existence of more subwavelength resonances and more subwavelength solutions than in the linear case. This article is concluded with Section 6 discussing final conclusions and an outlook for further research.
2 Prerequisites and notation
In what follows, $\operatorname{{\mathcal{D}}}\subset\mathbb{R}^{3}$ is a smooth bounded domain with connected components $B_{1},\ldots,B_{N}$ that have connected boundaries. Their characteristic functions are denoted by $1_{B_{j}}$ and are given by
$1_{B_{j}}(x)=\begin{cases}1\quad\text{if }x\in B_{j},\\ 0\quad\text{otherwise.}\end{cases}$
For $k\in\mathbb{C}$, the single layer potential $\mathcal{S}^{k}_{\operatorname{{\mathcal{D}}}}$ and the adjoint of the Neumann-Poincaré operator $\mathcal{K}^{k,*}_{\operatorname{{\mathcal{D}}}}$ are given for $\phi\in H^{-\frac{1}{2}}(\partial\operatorname{{\mathcal{D}}})$ by
$\mathcal{S}^{k}_{\operatorname{{\mathcal{D}}}}[\phi](x)$ $:=\int_{\partial\operatorname{{\mathcal{D}}}}G^{k}(x-y)\phi(y)\operatorname{d\sigma}(y),\quad x\in\partial\operatorname{{\mathcal{D}}},$
$\mathcal{K}^{k,*}_{\operatorname{{\mathcal{D}}}}[\phi](x)$ $:=\int_{\partial\operatorname{{\mathcal{D}}}}\frac{\partial G^{k}(x-y)}{\partial\nu(x)}\phi(y)\operatorname{d\sigma}(y),\quad x\in\partial\operatorname{{\mathcal{D}}},$
where $\nu(x)$ denotes the outward pointing normal vector to $\partial\operatorname{{\mathcal{D}}}$ at $x\in\partial\operatorname{{\mathcal{D}}}$ and the Green’s function $G^{k}$ of the Helmholtz equation with radiation condition is given by
$G^{k}(x-y):=-\frac{\exp(ik\lvert x-y\rvert)}{4\pi\lvert x-y\rvert}.$
Let $\phi\in H^{\frac{1}{2}}(\partial\operatorname{{\mathcal{D}}})$. Recall that the following system:
$\left\{\begin{array}[]{l}\Delta u+k^{2}u=~{}~{}0\quad\text{ in }\mathbb{R}^{3}\setminus\overline{\operatorname{{\mathcal{D}}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr u=~{}~{}\phi\quad\text{ on }\partial\operatorname{{\mathcal{D}}},\end{array}\right.$
subject to an outgoing radiation condition, has a unique solution. In that case the Dirichlet-to-Neumann operator $\mathcal{T}^{k}_{\operatorname{{\mathcal{D}}}}[\phi]$ is defined as
$\mathcal{T}^{k}_{\operatorname{{\mathcal{D}}}}[\phi](x):=\frac{\partial u}{\partial\nu}(x),\quad x\in\partial\operatorname{{\mathcal{D}}}.$
The Dirichlet-to-Neumann operator can be expanded into a type of generalized Taylor series in the space of bounded linear operators $\mathcal{L}(H^{\frac{1}{2}}(\partial\operatorname{{\mathcal{D}}}),H^{-\frac{1}{2}}(\partial\operatorname{{\mathcal{D}}}))$ for $k$ in a neighborhood of $0\in\mathbb{C}$. The following functions will be needed for the linear term of the Taylor series of $\mathcal{T}^{k}_{\operatorname{{\mathcal{D}}}}$.
Definition 2.1.
Let $j\in\{1,\ldots,N\}$. Define the following functions $\Phi_{j}:\operatorname{{\mathbb{R}^{3}}}3\setminus\overline{\operatorname{{\mathcal{D}}}}\rightarrow\mathbb{C}$ as the solution to the following system of equations:
$\left\{\begin{array}[]{rll}\Delta\Phi_{j}&=~{}~{}0&\text{in }\operatorname{{\mathbb{R}^{3}}}3\setminus\overline{\operatorname{{\mathcal{D}}}},\\ \Phi_{j}|_{\partial\operatorname{{\mathcal{D}}}}&=~{}~{}1_{\partial B_{j}}&\text{ on }\partial\operatorname{{\mathcal{D}}}.\end{array}\right.$
With these objects it is now possible to formulate the following lemma. The next results can be deduced from [feppon2024subwavelength, Proposition 3.1] and properties of the single layer potential which, e.g., can be found in [photonic].
Lemma 2.2.
Let $k\in\mathbb{C}$ and suppose that $k^{2}$ is not a Dirichlet eigenvalue for $-\Delta$ in $\operatorname{{\mathcal{D}}}$. Then, the following statements hold:
  1. (i)
    The single layer potential $\mathcal{S}^{k}_{\operatorname{{\mathcal{D}}}}:H^{-\frac{1}{2}}(\partial\operatorname{{\mathcal{D}}})\rightarrow H^{\frac{1}{2}}(\partial\operatorname{{\mathcal{D}}})$ is invertible;
  2. (ii)
    The Dirichlet-to-Neumann operator $\mathcal{T}^{k}_{\operatorname{{\mathcal{D}}}}:H^{\frac{1}{2}}(\partial\operatorname{{\mathcal{D}}})\rightarrow H^{-\frac{1}{2}}(\partial\operatorname{{\mathcal{D}}})$ is given by
    $\mathcal{T}^{k}_{\operatorname{{\mathcal{D}}}}[\phi]=\left(\frac{1}{2}\operatorname{Id}+\mathcal{K}^{k,*}_{\operatorname{{\mathcal{D}}}}\right)\left(\mathcal{S}_{\operatorname{{\mathcal{D}}}}^{k}\right)^{-1}[\phi],$
    where $\operatorname{Id}$ denotes the identity operator;
  3. (iii)
    For $k\rightarrow 0\in\mathbb{C}$, the Dirichlet-to-Neumann operator and the single layer potential can be expanded as a convergent power series in the spaces of bounded linear operators $\mathcal{L}(H^{\frac{1}{2}}(\partial\operatorname{{\mathcal{D}}}),H^{-\frac{1}{2}}(\partial\operatorname{{\mathcal{D}}}))$ and $\mathcal{L}(H^{-\frac{1}{2}}(\partial\operatorname{{\mathcal{D}}}),H^{\frac{1}{2}}(\partial\operatorname{{\mathcal{D}}}))$, respectively. The following identities hold:
    $\mathcal{T}^{k}_{\operatorname{{\mathcal{D}}}}=\sum_{n\geq 0}k^{n}\mathcal{T}_{n},\quad\mathcal{S}^{k}_{\operatorname{{\mathcal{D}}}}=\sum_{n\geq 0}k^{n}\mathcal{S}_{n},$
    where
    $\mathcal{T}_{0}=\mathcal{T}^{0}_{\operatorname{{\mathcal{D}}}},\quad\quad\mathcal{S}_{0}=\mathcal{S}^{0}_{\operatorname{{\mathcal{D}}}}\quad\text{ and }\quad\mathcal{T}_{1}[\phi]=-\mathcal{T}^{0}_{\operatorname{{\mathcal{D}}}}\circ\mathcal{S}_{1}\circ(\mathcal{S}^{0}_{\operatorname{{\mathcal{D}}}})^{-1}.$
For this work, the most important part of this Lemma 2.2 will be item (iii). The formula provided for $\mathcal{T}_{0}$ and $\mathcal{T}_{1}$ will be needed for determining the subwavelength resonances asymptotically in terms of $\delta$ in Subsection 3.2 and Subsection 4.2.
In the introduction (Section 1), the fundamental theorem of subwavelength physics was presented in Theorem 1.2. The statement is that subwavelength resonances (in the linear setting) correspond to the eigenvalues of a matrix $\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})$ called the generalized capacitance matrix. The precise definition of the capacitance matrix is as follows [ammari2017plasmonic].
Definition 2.3 (Capacitance matrix).
Let $\operatorname{{\mathcal{D}}}\subset\mathbb{R}^{3}$ a smooth bounded domain with connected components $B_{1},\ldots,B_{N}$ that are well separated and have connected boundaries. The capacitance matrix $\operatorname{cap}(\operatorname{{\mathcal{D}}})\in\mathbb{R}^{N\times N}$ associated to $\operatorname{{\mathcal{D}}}$ is defined as
$\operatorname{cap}(\operatorname{{\mathcal{D}}})_{l,j}=-\int_{\partial B_{l}}\frac{\partial\Phi_{j}}{\partial\nu}\operatorname{d\sigma},$
where $\Phi_{j}$ is as in Definition 2.1. Let $\lvert B_{j}\rvert$ denote the volume of the $B_{j}\subset\operatorname{{\mathbb{R}^{3}}}3$ and ${c_{r}^{j}}$ the material parameter in $B_{j}$, for $j=1,\ldots,N$, and let $V^{vol}$ be the diagonal matrix with the diagonal entries defined as
$(V^{vol})_{j,j}=\frac{{c_{r}^{j}}^{2}}{\lvert B_{j}\rvert}.$
The generalized capacitance matrix $\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})\in\mathbb{C}^{N\times N}$ associated to $\operatorname{{\mathcal{D}}}$ and the material parameter ${c_{r}}$ is defined as
$\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})=V^{vol}\operatorname{cap}(\operatorname{{\mathcal{D}}}).$
Note that if ${c_{r}^{j}}={c_{r}}$ for all $j$, then
$\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})={c_{r}}^{2}V^{vol}\operatorname{cap}(\operatorname{{\mathcal{D}}}).$
In Section 3, we will reproduce Theorem 1.2 with the techniques of Dirichlet-to-Neumann operator and a variational approach.
3 A Dirichlet-to-Neumann approach to subwavelength resonances
This section is devoted to reproducing the fundamental theorem of subwavelength physics (Theorem 1.2) with a novel technique, first used for this purpose in [feppon2024subwavelength]. The original derivation was based on layer potential techniques and the asymptotic Gohberg-Sigal theory (as it can be found for example in [photonic]). In this work, we take a different approach, which is based on Dirichlet-to-Neumann operators and variational arguments. To this end, we pose the following proposition which reformulates the system of equations (2) in terms of the Dirichlet-to-Neumann operator as in Section 2 and as a variational problem in $H^{1}(\operatorname{{\mathcal{D}}})$.
Proposition 3.1.
Let $u\in H^{1}(\operatorname{{\mathcal{D}}})$. Then the following statements are equivalent:
  1. (i)
    The function $u$ solves the system (2) for some outgoing $u|_{\mathbb{R}^{3}\setminus\overline{\operatorname{{\mathcal{D}}}}}\in H^{1}_{loc}(\mathbb{R}^{3}\setminus\overline{\operatorname{{\mathcal{D}}}})$;
  2. (ii)
    The function $u$ satisfies
    $\left\{\begin{array}[]{rll}\displaystyle\Delta u+\frac{\omega^{2}}{{c_{r}}^{2}}u&=0&\text{in }\operatorname{{\mathcal{D}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\frac{\partial u}{\partial\nu}&=\displaystyle\delta\left(\mathcal{T}_{\operatorname{{\mathcal{D}}}}^{\frac{\omega}{{c_{0}}}}[u-u_{in}|_{+}]+\left.\frac{\partial u_{in}}{\partial\nu}\right|_{+}\right)&\text{on }\partial\operatorname{{\mathcal{D}}};\end{array}\right.$
  3. (iii)
    The function $u$ solves the variational problem
    $\begin{split}\int_{\operatorname{{\mathcal{D}}}}\nabla u\cdot\nabla\overline{v}\operatorname{dvol}-\frac{\omega^{2}}{{c_{r}}^{2}}\int_{\operatorname{{\mathcal{D}}}}u\overline{v}\operatorname{dvol}-\delta\int_{\partial\operatorname{{\mathcal{D}}}}\mathcal{T}_{\operatorname{{\mathcal{D}}}}^{\frac{\omega}{{c_{0}}}}[u]\overline{v}\operatorname{d\sigma}=\delta\int_{\partial\operatorname{{\mathcal{D}}}}\left(\left.\frac{\partial u_{in}}{\partial\nu}\right|_{+}-\mathcal{T}_{\operatorname{{\mathcal{D}}}}^{\frac{\omega}{{c_{0}}}}[u_{in}|_{+}]\right)\overline{v}\operatorname{d\sigma},\\ \quad\forall v\in H^{1}(\operatorname{{\mathcal{D}}}).\end{split}$ (7)
Proof.
Assume that $u$ solves the system (2) for some outgoing $u|_{\mathbb{R}^{3}\setminus\overline{\operatorname{{\mathcal{D}}}}}\in H^{1}_{loc}(\mathbb{R}^{3}\setminus\overline{\operatorname{{\mathcal{D}}}})$. Then $\left.\left(u|_{\mathbb{R}^{3}\setminus\overline{\operatorname{{\mathcal{D}}}}}\right)\right|_{+}=u|_{-}-u_{in}$ and $u_{\mathbb{R}^{3}\setminus\overline{\operatorname{{\mathcal{D}}}}}$ is outgoing and solves the exterior Helmholtz problem. Thus,
$\left.\frac{\partial u|_{\mathbb{R}^{3}\setminus\overline{\operatorname{{\mathcal{D}}}}}}{\partial\nu}\right|_{+}=\mathcal{T}^{\frac{\omega}{{c_{0}}}}_{\operatorname{{\mathcal{D}}}}\left[\left.\left(u|_{\mathbb{R}^{3}\setminus\overline{\operatorname{{\mathcal{D}}}}}\right)\right|_{+}\right]=\mathcal{T}^{\frac{\omega}{{c_{0}}}}_{\operatorname{{\mathcal{D}}}}[u|_{-}-u_{in}].$
Hence,
$\left.\frac{\partial u}{\partial\nu}\right|_{-}=\delta\left(\mathcal{T}_{\operatorname{{\mathcal{D}}}}^{\frac{\omega}{{c_{0}}}}[u|_{-}-u_{in}|_{+}]+\left.\frac{\partial u_{in}}{\partial\nu}\right|_{+}\right)$
and (ii) follows. Clearly, if $u$ satisfies (ii), then there exists an outgoing $u|_{\mathbb{R}^{3}\setminus\operatorname{{\mathcal{D}}}}\in H^{1}_{loc}(\mathbb{R}^{3}\setminus\overline{\operatorname{{\mathcal{D}}}})$ such that
$\left.\left(u|_{\mathbb{R}^{3}\setminus\operatorname{{\mathcal{D}}}}\right)\right|_{+}=u|_{-}-u_{in}|_{+}\quad\text{and}\quad\left.\frac{\partial u|_{\mathbb{R}^{3}\setminus\overline{\operatorname{{\mathcal{D}}}}}}{\partial\nu}\right|_{+}=\mathcal{T}^{\frac{\omega}{{c_{0}}}}_{\operatorname{{\mathcal{D}}}}\left[u|_{-}-u_{in}|_{+}\right].$
Suppose that $u\in H^{1}(\operatorname{{\mathcal{D}}})$ satisfies (ii) and let $v\in H^{1}(\operatorname{{\mathcal{D}}})$ be arbitrary, multiplying the first equation in (ii) with $\overline{v}$ and integrating by parts over $\operatorname{{\mathcal{D}}}$ gives
$\int_{\operatorname{{\mathcal{D}}}}\left(\Delta u+\frac{\omega^{2}}{{c_{r}}^{2}}u\right)\overline{v}\operatorname{dvol}=\int_{\operatorname{{\mathcal{D}}}}-\nabla u\cdot\nabla\overline{v}+\frac{\omega^{2}}{{c_{r}}^{2}}u\overline{v}\operatorname{dvol}+\int_{\partial\operatorname{{\mathcal{D}}}}\frac{\partial u}{\partial\nu}\overline{v}\operatorname{d\sigma}.$
Inserting the second equation of (ii) for ${\partial u}/{\partial\nu}$ one obtains that $u$ solves the variational problem in (iii). Suppose that $u$ solves the variational problem in (iii). In order to recover that $u$ satisfies (ii), one first considers $v\in H^{1}_{0}(\operatorname{{\mathcal{D}}})$ and derives (using integration by parts), that $u$ satisfies the Helmholtz equation in $\operatorname{{\mathcal{D}}}$, considering then a general $v\in H^{1}(\operatorname{{\mathcal{D}}})$ allows to conclude the equality for the boundary condition in (ii). ∎
An immediate consequence of Proposition 3.1 is the following lemma that characterizes variationally the resonances of the system (2).
Corollary 3.2.
Let $\omega(\delta)\in\mathbb{C}$. Then, $\omega(\delta)$ is a resonance of (2) if and only if there exists a non-zero $u\in H^{1}(\operatorname{{\mathcal{D}}})$ such that
$\int_{\operatorname{{\mathcal{D}}}}\nabla u\cdot\nabla\overline{v}\operatorname{dvol}-\frac{\omega(\delta)^{2}}{{c_{r}}^{2}}\int_{\operatorname{{\mathcal{D}}}}u\overline{v}\operatorname{dvol}-\delta\int_{\partial\operatorname{{\mathcal{D}}}}\mathcal{T}^{\frac{\omega(\delta)}{{c_{0}}}}[u]\overline{v}\operatorname{d\sigma}=0,\quad\forall v\in H^{1}(\operatorname{{\mathcal{D}}}).$
In other words, $\omega(\delta)$ is a resonance whenever the variational problem (7) does not admit a unique solution.
In the following, we will restrict our study to the subwavelength regime. That is, we will consider $\delta$ in a neighborhood of $0\in\mathbb{R}$ and $\omega$ in a neighborhood of $0\in\mathbb{C}$. We will enlarge the variational framework given in [feppon2024subwavelength] to the case of a disconnected resonator domain $\operatorname{{\mathcal{D}}}$ and characterize the subwavelength resonances in that case.
The idea is as follows. The bilinear form
$(u,v)\longmapsto\int_{\operatorname{{\mathcal{D}}}}\nabla u\cdot\nabla\overline{v}\operatorname{dvol}-\frac{\omega^{2}}{{c_{r}}^{2}}\int_{\operatorname{{\mathcal{D}}}}u\overline{v}\operatorname{dvol}-\delta\int_{\partial\operatorname{{\mathcal{D}}}}\mathcal{T}_{\operatorname{{\mathcal{D}}}}^{\frac{\omega}{{c_{0}}}}[u]\overline{v}\operatorname{d\sigma},$
is degenerate for a continuity of $(\omega,\delta)$ in a neighborhood of $0\in\mathbb{C}\times\mathbb{R}$, namely the subwavelength resonances $(\omega(\delta),\delta)$. While these characteristic points are the objects of interest, it is unfortunate that the problem is in some sense indirect. This complicates the analysis. The great advantage of the method presented in [feppon2024subwavelength] is that it translates the indirect problem of finding characteristic points to a direct problem, where one solves a well-posed variational problem. The solution $u_{\omega,\delta}$ of that problem is then used to determine whether $(\omega,\delta)$ is a resonance. Subsection 3.1 will treat the first step of developing the well-posed variational problem. In Subsection 3.2, the solution of that problem will be computed and a characterization of subwavelength resonances will be derived.
3.1 Variational formulation
In this section, we will develop a variational characterization of resonances of the linear problem (2). To this end, we pose the following definition.
Definition 3.3.
Define the bilinear form $a_{\omega,\delta}:H^{1}(\operatorname{{\mathcal{D}}})\times H^{1}(\operatorname{{\mathcal{D}}})\rightarrow\mathbb{C}$ associated to the smooth domain $\operatorname{{\mathcal{D}}}$ with well-separated connected components $B_{1},\ldots,B_{N}$ that have a connected boundaries,
$a_{\omega,\delta}(u,v):=\int_{\operatorname{{\mathcal{D}}}}\nabla u\cdot\nabla\overline{v}\operatorname{dvol}+\sum_{j=1}^{N}\int_{B_{j}}u\operatorname{dvol}\int_{B_{j}}\overline{v}\operatorname{dvol}-\frac{\omega^{2}}{{c_{r}}^{2}}\int_{\operatorname{{\mathcal{D}}}}u\overline{v}\operatorname{dvol}-\delta\int_{\partial\operatorname{{\mathcal{D}}}}\mathcal{T}_{\operatorname{{\mathcal{D}}}}^{\frac{\omega}{{c_{0}}}}[u]\overline{v}\operatorname{d\sigma}.$ (8)
The resonances of the linear Helmholtz problem (2) will be characterized in terms of the bilinear form $a_{\omega,\delta}$ in Theorem 3.5. In the proof of Theorem 3.5, the following lemma from [feppon2024subwavelength] will be useful.
Lemma 3.4.
There exists a neighborhood $U$ of $0\in\mathbb{R}$ and a neighborhood $V$ of $0\in\mathbb{C}$ such that for all $\delta\in U$ and for all $\omega\in V$ and for any $f\in H^{1}(\operatorname{{\mathcal{D}}})^{*}$, the dual of $H^{1}(\operatorname{{\mathcal{D}}})$, the following variational problem has a unique solution $u_{f}(\omega,\delta)$
$a_{\omega,\delta}(u_{f}(\omega,\delta),v)=\langle f,v\rangle_{H^{1}(\operatorname{{\mathcal{D}}})^{*},H^{1}(\operatorname{{\mathcal{D}}})},\quad\forall\,v\in H^{1}(\operatorname{{\mathcal{D}}}).$ (9)
Theorem 3.5.
Let $\omega\in\mathbb{C}$ and $\delta\in\mathbb{R}$ belong to a neighborhood of zero. For any $f\in H^{1}(\operatorname{{\mathcal{D}}})^{*}$, the variational problem:
[…]
admits a unique solution $u$ if and only if
$\displaystyle\operatorname{Id}-\left(\int_{B_{i}}u_{1,j}\operatorname{dvol}\right)_{i,j}$ (11)
is invertible, where $u_{1,j}\in H^{1}(\operatorname{{\mathcal{D}}})$ is the unique solution to the variational problem
$a_{\omega,\delta}(u_{1,j}(\omega,\delta),v)=\langle 1_{B_{j}},v\rangle_{H^{1}(\operatorname{{\mathcal{D}}})^{*},H^{1}(\operatorname{{\mathcal{D}}})}=\int_{B_{j}}\overline{v}\operatorname{dvol},\quad\forall\,v\in H^{1}(\operatorname{{\mathcal{D}}}).$ (12)
Proof.
Clearly, (10) is equivalent to
$a_{\omega,\delta}(u,v)-\sum_{j=1}^{N}\int_{B_{j}}u\operatorname{dvol}\int_{B_{j}}\overline{v}\operatorname{dvol}=\langle f,v\rangle_{H^{1}(\operatorname{{\mathcal{D}}})^{*},H^{1}(\operatorname{{\mathcal{D}}})}$
$\iff a_{\omega,\delta}(u,v)=a_{\omega,\delta}(u_{f}(\omega,\delta),v)+\sum_{j=1}^{N}\left(\int_{B_{j}}u\operatorname{dvol}\right)a_{\omega,\delta}(u_{1,j}(\omega,\delta),v),$
which implies
$u=u_{f}(\omega,\delta)+\sum_{j=1}^{N}\left(\int_{B_{j}}u\operatorname{dvol}\right)u_{1,j}(\omega,\delta).$
Let $\delta_{l,j}$ denote the Kronecker delta. Integrating the above equation on $B_{l}$ gives the following system of equations:
$\begin{pmatrix}\displaystyle\int_{B_{l}}u_{f}(\omega,\delta)\operatorname{dvol}\end{pmatrix}_{l}=\sum_{j=1}^{N}\left(\delta_{l,j}-\begin{pmatrix}\int_{B_{l}}u_{1,j}(\omega,\delta)\operatorname{dvol}\end{pmatrix}_{l,j}\right)\begin{pmatrix}\int_{B_{j}}u\operatorname{dvol}\end{pmatrix}_{j},$
which has a unique solution if and only if the matrix
$\operatorname{Id}-\begin{pmatrix}\int_{B_{l}}u_{1,j}(\omega,\delta)\operatorname{dvol}\end{pmatrix}_{l,j}$
is invertible. ∎
This theorem allows to characterize the resonances as precisely those $\omega(\delta)\in\mathbb{C}$ such that the matrix in (11) is degenerate. We have the following corollary.
Corollary 3.6.
Let $\omega\in\mathbb{C}$ and $\delta\in\mathbb{R}$ and suppose that $\omega$ and $\delta$ are sufficiently small. Then, $\omega$ is a resonance of the system (2) if and only if (11) is degenerate.
3.2 Characterization of the resonances
In this Subsection the solution of the variational problem (12) will be determined for small $\omega\in\mathbb{C}$, $\delta\in\mathbb{R}$ and then, by using Corollary 3.6, the subwavelength resonances will be computed. To this end, we consider the PDE formulation of the variational problem (12)
$\left\{\begin{array}[]{rl}\displaystyle-\Delta u_{1,j}+\sum_{l=1}^{N}\left(\int_{B_{l}}u_{1,j}\operatorname{dvol}\right)1_{B_{l}}&=~{}~{}\displaystyle\frac{\omega^{2}}{{c_{r}}^{2}}u_{1,j}+1_{B_{j}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\frac{\partial u_{1,j}}{\partial\nu}&=~{}~{}\delta\mathcal{T}^{\frac{\omega}{{c_{0}}}}[u_{1,j}].\end{array}\right.$ (13)
We will furthermore need the following proposition.
Proposition 3.7.
The following equality holds:
$\mathcal{T}_{1}[1_{B_{j}}]=-\frac{i}{4\pi}\sum_{l=1}^{N}\operatorname{cap}(\operatorname{{\mathcal{D}}})_{l,j}\sum_{l=1}^{N}\frac{\partial\Phi_{l}}{\partial\nu}.$
Proof.
Using item (iii) in Lemma 2.2 together with Definition 2.3 and the fact from [feppon2022modal, Proposition 12] that
$\mathcal{S}_{1}[\phi]=-\frac{i}{4\pi}\int_{\partial\operatorname{{\mathcal{D}}}}\phi\operatorname{d\sigma},$
we obtain that
$\mathcal{T}_{1}[1_{B_{j}}]=-\mathcal{T}^{0}_{\operatorname{{\mathcal{D}}}}\circ\mathcal{S}_{1}((\mathcal{S}^{0}_{\operatorname{{\mathcal{D}}}})^{-1}[1_{B_{j}}])=\frac{i}{4\pi}\mathcal{T}^{0}_{\operatorname{{\mathcal{D}}}}[1_{\operatorname{{\mathcal{D}}}}]\int_{\operatorname{{\mathcal{D}}}}\frac{\partial\Phi_{j}}{\partial\nu}\operatorname{dvol}=-\sum_{l=1}^{N}\frac{\partial\Phi_{l}}{\partial\nu}\sum_{l=1}^{N}\operatorname{cap}(\operatorname{{\mathcal{D}}})_{l,j}.$
With Proposition 3.7, it is now possible to compute $u_{1,j}(\omega,\delta)$ from Theorem 3.5 asymptotically as $\omega,\delta\rightarrow 0$.
Proposition 3.8.
The solution $u_{1,j}(\omega,\delta)$ to (12) has the following asymptotic behavior as $\omega,\delta\rightarrow 0$:
$u_{1,j}(\omega,\delta)=\frac{1}{\lvert B_{j}\rvert}1_{B_{j}}+\omega^{2}\frac{1}{{c_{r}}^{2}\lvert B_{j}\rvert^{2}}1_{B_{j}}+\delta\left(-\sum_{l=1}^{N}\frac{\operatorname{cap}(\operatorname{{\mathcal{D}}})_{l,j}}{\lvert B_{j}\rvert\lvert B_{l}\rvert^{2}}1_{B_{l}}+\tilde{u}^{j}_{0,1}\right)$
$+\omega\delta\left(\sum_{l=1}^{N}\left(\frac{i}{4\pi\lvert B_{j}\rvert\lvert B_{l}\rvert^{2}{c_{0}}}(\operatorname{cap}(\operatorname{{\mathcal{D}}})J\operatorname{cap}(\operatorname{{\mathcal{D}}}))_{l,j}\right)1_{B_{l}}+\tilde{u}^{j}_{1,1}\right)+O((\omega^{2}+\delta)^{2}),$
where $J$ denotes the $(N\times N)$-matrix of ones.
Proof.
We make the Ansatz
$u_{1,j}(\omega,\delta)=\sum_{p,k\geq 0}\omega^{p}\delta^{k}u_{p,k},$
inserting this into (13) and identifying powers of $\delta$ and $\omega$, we obtain the following cascade of equations characterizing the functions $(u_{p,k})_{p,k\geq 0}$:
$-\Delta u_{p,k}+\sum_{i=1}^{N}\left(\int_{B_{i}}u_{p,k}\operatorname{dvol}\right)1_{B_{i}}$ $=\frac{1}{{c_{r}}^{2}}u_{p-2,k}+1_{B_{j}}\delta_{p=0}\delta_{k=0},$
$\frac{\partial u_{p,k}}{\partial\nu}$ $=\sum_{n=0}^{p}\frac{1}{{c_{0}}^{n}}\mathcal{T}_{n}[u_{p-n,k-1}].$
First, we consider $p\geq 0$ and $k=0$
$-\Delta u_{p,0}+\sum_{i=1}^{N}\left(\int_{B_{i}}u_{p,0}\operatorname{dvol}\right)1_{B_{i}}$ $=\frac{1}{{c_{r}}^{2}}u_{p-2,0}+1_{B_{j}}\delta_{p=0}\delta_{0=0},$
$\frac{\partial u_{p,0}}{\partial\nu}$ $=0.$
One can prove inductively that
$u_{2p,0}=\frac{1}{c_{r}^{2p}\lvert B_{j}\rvert^{p+1}}1_{B_{j}},\quad u_{2p+1,0}=0,\quad\text{ for }\quad p\geq 0.$
Let $p=0,k=1$. Then $u_{0,1}$ satisfies the following system of equations:
$-\Delta u_{0,1}+\sum_{i=1}^{N}\left(\int_{B_{i}}u_{0,1}\operatorname{dvol}\right)1_{B_{i}}$ $=0,$
$\frac{\partial u_{0,1}}{\partial\nu}$ $=\mathcal{T}_{0}[u_{0,0}].$
The integral of $u_{0,1}$ on $B_{l}$ is thus given by
$\lvert B_{l}\rvert\int_{B_{l}}u_{0,1}\operatorname{dvol}=\int_{\partial B_{l}}\frac{\partial u_{0,1}}{\partial\nu}\operatorname{d\sigma}=\int_{\partial B_{l}}\mathcal{T}^{0}[u_{0,0}]\operatorname{d\sigma}=\frac{1}{\lvert B_{j}\rvert}\int_{\partial B_{l}}\frac{\partial\Phi_{j}}{\partial\nu}\operatorname{d\sigma}=-\frac{1}{\lvert B_{j}\rvert}\operatorname{cap}(\operatorname{{\mathcal{D}}})_{l,j}.$
It follows that $u_{0,1}$ is given by
$u_{0,1}=-\sum_{l=1}^{N}\frac{1}{\lvert B_{j}\rvert\lvert B_{l}\rvert^{2}}\operatorname{cap}(\operatorname{{\mathcal{D}}})_{l,j}1_{B_{l}}+\tilde{u}_{0,1}^{j},$
where $\tilde{u}_{0,1}^{j}$ satisfies
$\left\{\begin{array}[]{rll}\Delta\tilde{u}_{0,1}^{j}&=~{}~{}0&\text{in }\operatorname{{\mathcal{D}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\frac{\partial\tilde{u}_{0,1}^{j}}{\partial\nu}&=~{}~{}\displaystyle\frac{1}{\lvert B_{j}\rvert}\frac{\partial\Phi_{j}}{\partial\nu}&\text{on }\partial\operatorname{{\mathcal{D}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\int_{B_{l}}\tilde{u}_{0,1}^{j}\operatorname{dvol}&=~{}~{}0,&\forall~{}l\in\{1,\ldots,N\}.\end{array}\right.$
For $p,k=1,$ one obtains that
$-\Delta u_{1,1}+\sum_{i=1}^{N}\left(\int_{B_{i}}u_{1,1}\operatorname{dvol}\right)1_{B_{i}}$ $=0,$
$\frac{\partial u_{1,1}}{\partial\nu}$ $=\mathcal{T}_{0}[u_{1,0}]+\frac{1}{{c_{0}}}\mathcal{T}_{1}[u_{0,0}]=\frac{1}{{c_{0}}}\mathcal{T}_{1}[u_{0,0}].$
Using Proposition 3.7 yields
$\frac{1}{{c_{0}}}\mathcal{T}_{1}[u_{0,0}]=\frac{1}{{c_{0}}\lvert B_{j}\rvert}\mathcal{T}_{1}[1_{B_{j}}]$ $=-\frac{i}{4\pi{c_{0}}\lvert B_{j}\rvert}\sum_{l=1}^{N}\operatorname{cap}(\operatorname{{\mathcal{D}}})_{l,j}\sum_{l=1}^{N}\frac{\partial\Phi_{l}}{\partial\nu}.$
Therefore, it follows that
$\lvert B_{m}\rvert\int_{B_{m}}u_{1,1}\operatorname{dvol}=\frac{-i}{4\pi{c_{0}}\lvert B_{j}\rvert}\sum_{l=1}^{N}\operatorname{cap}(\operatorname{{\mathcal{D}}})_{l,j}\int_{\partial B_{m}}\sum_{l=1}^{N}\frac{\partial\Phi_{l}}{\partial\nu}\operatorname{d\sigma}$
$=\frac{i}{4\pi{c_{0}}\lvert B_{j}\rvert}\sum_{l=1}^{N}\operatorname{cap}(\operatorname{{\mathcal{D}}})_{l,j}\sum_{l=1}^{N}\operatorname{cap}(\operatorname{{\mathcal{D}}})_{m,l}$
$=\frac{i}{4\pi{c_{0}}\lvert B_{j}\rvert}(\operatorname{cap}(\operatorname{{\mathcal{D}}})J\operatorname{cap}(\operatorname{{\mathcal{D}}}))_{m,j},$
and we can decompose $u_{1,1}$ as
$u_{1,1}=\sum_{m=1}^{N}\frac{i}{4\pi{c_{0}}\lvert B_{j}\rvert\lvert B_{m}\rvert^{2}}(\operatorname{cap}(\operatorname{{\mathcal{D}}})J\operatorname{cap}(\operatorname{{\mathcal{D}}}))_{m,j}1_{B_{m}}+\tilde{u}_{1,1}^{j},$
where $\tilde{u}_{1,1}^{j}$ is the solution to
$\left\{\begin{array}[]{rll}\Delta\tilde{u}_{1,1}^{j}&=~{}~{}0&\text{in }\operatorname{{\mathcal{D}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\frac{\partial\tilde{u}_{1,1}^{j}}{\partial\nu}&=~{}~{}\displaystyle-\frac{i}{4\pi{c_{0}}\lvert B_{j}\rvert}\sum_{l=1}^{N}\operatorname{cap}(\operatorname{{\mathcal{D}}})_{l,j}\sum_{l=1}^{N}\frac{\partial\Phi_{l}}{\partial\nu}&\text{on }\partial\operatorname{{\mathcal{D}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\int_{B_{l}}\tilde{u}_{1,1}^{j}\operatorname{dvol}&=~{}~{}0,&\forall~{}l\in\{1,\ldots,N\}.\end{array}\right.$
Applying Corollary 3.6 one can reproduce the fundamental theorem of subwavelength physics, Theorem 1.2.
Theorem 3.9 (Fundamental theorem of subwavelength physics).
Let $\{q_{0},v_{2},\ldots,v_{N}\}$ be a basis of eigenvectors of $\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})$ and denote by $\Pi:\mathbb{C}^{N}\rightarrow\mathbb{C}q_{0}$ the projection along $\operatorname{span}(v_{2},\ldots,v_{N})$ onto $\mathbb{C}q_{0}$. Then the subwavelength resonances $\omega(\delta)=\omega_{0}\sqrt{\delta}+\omega_{1}\delta+O(\delta^{\frac{3}{2}})$ of (2) associated to $q_{0}$ satisfies the following asymptotics:
$\omega_{0}\text{ is an eigenvector of }\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})\text{ with eigenvector }q_{0},$
$\omega_{1}=-\frac{i}{8\pi{c_{0}}}\frac{q_{0}\cdot\Pi\left[\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})J\operatorname{cap}(\operatorname{{\mathcal{D}}})q_{0}\right]}{\lVert q_{0}\rVert^{2}}.$
Proof.
By Theorem 3.5, the subwavelength resonances of (2) are given by those $\omega(\delta)$ such that (11) is not invertible. Inserting the results of Proposition 3.8 into (11) implies that $\omega(\delta)$ is a subwavelength resonance whenever there exists $q(\delta)\in\mathbb{C}^{N}\setminus\{0\}$ such that
$\left(\omega(\delta)^{2}-\delta\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})-\omega(\delta)\delta\frac{i}{4\pi{c_{0}}}\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})J\operatorname{cap}(\operatorname{{\mathcal{D}}})\right)q(\delta)+O((\omega(\delta)^{2}+\delta)^{2})=0.$
Making the Ansatz $\omega(\delta)=\omega_{0}\sqrt{\delta}+\omega_{1}\delta+O(\delta^{\frac{3}{2}})$ and $q(\delta)=q_{0}+q_{1}\sqrt{\delta}+O(\delta)$ and inserting it into the above equation leads to the following cascade of equations:
$\left(\omega_{0}^{2}-\delta\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})\right)q_{0}=0,$
$\left(\omega_{0}^{2}-\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})\right)q_{1}+\omega_{0}\left(2\omega_{1}q_{0}-\frac{i}{4\pi{c_{0}}}\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})J\operatorname{cap}(\operatorname{{\mathcal{D}}})q_{0}\right)=0.$
It then follows that $\omega_{0}$ is an eigenvector of $\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})$ with an eigenvector $q_{0}$ and $\omega_{1}$ is given by
$\omega_{1}=\frac{q_{0}\cdot\Pi\left[\frac{i}{8\pi{c_{0}}}\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})J\operatorname{cap}(\operatorname{{\mathcal{D}}})q_{0}\right]}{\lVert q_{0}\rVert^{2}}.$
In the next Section, a fundamental theorem of nonlinear subwavelength physics will be derived.
4 Nonlinear subwavelength resonances
In this section, similar techniques to those in Section 3 are used to derive a fundamental theorem of nonlinear subwavelength physics, Theorem 4.5. The main difference is that the bilinear form $a_{\omega,\delta}$ in Theorem 3.5 will not be bilinear in this section, but will account for the nonlinearity $g(u)$ in (1). Consequently, in a formulation similar to (11), it will be important to account for the different amplitudes and the nonlinear mixing of the different $u_{1,j}$. Concretely, the condition (11-12) in Theorem 3.5 becomes a nonlinear system (15-16) in one function $u_{f}(\omega,\delta)\in H^{1}(\operatorname{{\mathcal{D}}})$ in Theorem 4.3.
The remaining steps of computing $u_{f}(\omega,\delta)$ asymptotically and characterizing the subwavelength resonances are then identical to the linear setting. The asymptotics of $u_{f}(\omega,\delta)$ can be found in Proposition 4.4 and the fundamental theorem of nonlinear subwavelength physics is given in Theorem 4.5.
It is important to note that the techniques in this section are not limited to $g(u)=\lvert\frac{\partial u}{\partial t}\rvert^{2}\frac{\partial u}{\partial t}$, but can be used for any $g(u)$ that satisfies that its Fourier transform is of the order $O(\omega^{2})$.
Similar to Proposition 3.1, the system (6) can equivalently be formulated with the Dirichlet-to-Neumann operator as follows:
$\left\{\begin{array}[]{rll}\displaystyle\Delta u+\frac{\omega^{2}}{{c_{r}}^{2}}u&=i\beta\lvert\omega\rvert^{2}\omega\lvert u\rvert^{2}u&\text{in }\operatorname{{\mathcal{D}}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\frac{\partial u}{\partial\nu}&=\delta\mathcal{T}_{\operatorname{{\mathcal{D}}}}^{\frac{\omega}{{c_{0}}}}[u]&\text{on }\partial\operatorname{{\mathcal{D}}},\end{array}\right.$
with $u:\operatorname{{\mathcal{D}}}\rightarrow\mathbb{C}$.
This is equivalent to the following variational problem: Find $u\in H^{1}(\operatorname{{\mathcal{D}}})$ such that
$\begin{split}\int_{\operatorname{{\mathcal{D}}}}\nabla u\cdot\nabla\overline{v}\operatorname{dvol}-\frac{\omega^{2}}{{c_{r}}^{2}}\int_{\operatorname{{\mathcal{D}}}}u\overline{v}\operatorname{dvol}-\int_{\operatorname{{\mathcal{D}}}}i\beta\lvert\omega\rvert^{2}\omega\lvert u\rvert^{2}u\overline{v}\operatorname{dvol}-\delta\int_{\partial\operatorname{{\mathcal{D}}}}\mathcal{T}_{\operatorname{{\mathcal{D}}}}^{\frac{\omega}{{c_{0}}}}[u]\overline{v}\operatorname{d\sigma}=0,\quad\quad\\ \quad\forall v\in H^{1}(\operatorname{{\mathcal{D}}}).\end{split}$
A resonance $\omega(\delta)$ is given whenever there exists $u_{\omega,\delta}\in H^{1}(\partial\operatorname{{\mathcal{D}}})\setminus\{0\}$ solving this variational problem.
Definition 4.1 (Nonlinear subwavelength resonance).
Let $\omega:(0,\epsilon)\rightarrow\mathbb{C}$ be a continuous function for some $\epsilon>0$. We say that $\omega(\delta)$ is a resonance of the system (6), if for every $\delta\in(0,\epsilon)$ the variational problem (4) has a non-zero solution $u_{\omega,\delta}\in H^{1}(\operatorname{{\mathcal{D}}})$.
A resonance $\omega(\delta)$ is said to be subwavelength if it also satisfies $\omega(\delta)\rightarrow 0\text{ as }\delta\rightarrow 0$.
4.1 Variational formulation
The following 2-form will play the role of $a_{\omega,\delta}$ in the nonlinear setting.
Definition 4.2.
Suppose that the smooth, bounded domain $\operatorname{{\mathcal{D}}}\in\operatorname{{\mathbb{R}^{3}}}3$ has well-separated connected components $B_{1},\ldots,B_{N}$ with connected boundaries. Define the $2$-form $\tilde{a}_{\omega,\delta}:H^{1}(\operatorname{{\mathcal{D}}})\times H^{1}(\operatorname{{\mathcal{D}}})\rightarrow\mathbb{C}$ associated $\operatorname{{\mathcal{D}}}$ by
$\begin{split}\tilde{a}_{\omega,\delta}(u,v):=\int_{\operatorname{{\mathcal{D}}}}\nabla u\cdot\nabla\overline{v}\operatorname{dvol}+\sum_{j=1}^{N}\int_{B_{j}}u\operatorname{dvol}\int_{B_{j}}\overline{v}\operatorname{dvol}-\frac{\omega^{2}}{{c_{r}}^{2}}\int_{\operatorname{{\mathcal{D}}}}u\overline{v}\operatorname{dvol}\quad\quad\\ -\int_{\operatorname{{\mathcal{D}}}}i\beta\lvert\omega\rvert^{2}\omega\lvert u\rvert^{2}u\overline{v}\operatorname{dvol}-\delta\int_{\partial\operatorname{{\mathcal{D}}}}\mathcal{T}_{\operatorname{{\mathcal{D}}}}^{\frac{\omega}{{c_{0}}}}[u]\overline{v}\operatorname{d\sigma}.\end{split}$ (14)
Recall here that $H^{1}(\operatorname{{\mathcal{D}}})\subset L^{4}(\operatorname{{\mathcal{D}}})$. With the form $\tilde{a}_{\omega,\delta}$ it is possible to characterize the resonances of the nonlinear system (6) in a similar fashion as in the linear case in Theorem 3.5 and Corollary 3.6. To this end, we will assume that there exists a neighborhood $U$ of $0\in\mathbb{R}$ and a neighborhood $V$ of $0\in\mathbb{C}$ such that for all $\delta\in U$ and for all $\omega\in V$ and for any $f\in H^{1}(\operatorname{{\mathcal{D}}})^{*}$ the following variational problem:
$\tilde{a}_{\omega,\delta}(u_{f}(\omega,\delta),v)=\langle f,v\rangle_{H^{1}(\operatorname{{\mathcal{D}}})^{*},H^{1}(\operatorname{{\mathcal{D}}})},\quad\forall v\in H^{1}(\operatorname{{\mathcal{D}}}),$
has a unique solution $u_{f}(\omega,\delta)$. With this property at hand, the following variational characterization of small resonances of the nonlinear system (6) is possible.
Theorem 4.3 (Characterization of resonances).
Let $\omega\in\mathbb{C}$ and $\delta\in\mathbb{R}$ belong to sufficiently small neighborhoods of zero. Then $\omega$ is a resonance of the system (6), if and only if there exist $a_{1},\ldots,a_{N}\in\mathbb{C}$ not all zero such that the solution $u_{f}$ to
$\tilde{a}_{\omega,\delta}(u_{f},v)=\sum_{i=1}^{N}a_{i}\int_{B_{i}}\overline{v}\operatorname{dvol},\quad\forall v\in H^{1}(\operatorname{{\mathcal{D}}}),$ (15)
satisfies
$\int_{B_{i}}u_{f}\operatorname{dvol}=a_{i},\quad\forall i\in\{a_{1},\ldots,a_{N}\}.$ (16)
Proof.
Assume that $\omega$ is a resonance of the system (6). Then, there exists $u_{f}\in H^{1}(\operatorname{{\mathcal{D}}})\setminus\{0\}$ such that $u_{f}$ satisfies (4). Set $a_{i}:=\int_{B_{i}}u_{f}\operatorname{dvol}$. Then, $u_{f}$ satisfies
$\tilde{a}_{\omega,\delta}(u_{f},v)=\sum_{i=1}^{N}\int_{B_{i}}u_{f}\operatorname{dvol}\int_{B_{i}}\overline{v}\operatorname{dvol}=\sum_{i=1}^{N}a_{i}\int_{B_{i}}\overline{v}\operatorname{dvol},\quad\forall v\in H^{1}(\operatorname{{\mathcal{D}}}),$
where the $a_{i}$ are not all zero. Conversely, assume there exist $a_{1},\ldots,a_{N}\in\mathbb{C}$ not all zero and $u_{f}\in H^{1}(\operatorname{{\mathcal{D}}})$ as in the theorem. Let $v\in H^{1}(\operatorname{{\mathcal{D}}})$, then using (15) and (16) one obtains
$\int_{\operatorname{{\mathcal{D}}}}\nabla u_{f}\cdot\nabla\overline{v}\operatorname{dvol}-\frac{\omega^{2}}{{c_{r}}^{2}}\int_{\operatorname{{\mathcal{D}}}}u_{f}\overline{v}\operatorname{dvol}-\int_{\operatorname{{\mathcal{D}}}}i\beta\lvert\omega\rvert^{2}\omega\lvert u_{f}\rvert^{2}u_{f}\overline{v}\operatorname{dvol}-\delta\int_{\partial\operatorname{{\mathcal{D}}}}\mathcal{T}_{\operatorname{{\mathcal{D}}}}^{\frac{\omega}{{c_{0}}}}[u_{f}]\overline{v}\operatorname{d\sigma}$
$=\tilde{a}_{\omega,\delta}(u_{f},v)-\sum_{i=1}^{N}\int_{B_{i}}u_{f}\operatorname{dvol}\int_{B_{i}}\overline{v}\operatorname{dvol}$
$=0.$
Hence, $u_{f}$ satisfies the variational problem (4), is not identically zero and thus $\omega$ is a resonance. ∎
4.2 Characterization of the resonances
In this subsection, we will determine the solution of the variational problem (15) and then derive asymptotics for subwavelength resonances of the nonlinear system (6). To this end, we consider the PDE formulation of the variational problem (15)
$\left\{\begin{array}[]{rl}\displaystyle-\Delta u_{f}+\sum_{j=1}^{N}a_{j}\left(\int_{B_{j}}u_{f}\operatorname{dvol}\right)1_{B_{j}}&=~{}~{}\displaystyle\frac{\omega^{2}}{{c_{r}}^{2}}u_{f}+i\beta\lvert\omega\rvert^{2}\omega\lvert u_{f}\rvert^{2}u_{f}+\sum_{j=1}^{N}a_{j}1_{B_{j}},\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\cr\displaystyle\frac{\partial u_{f}}{\partial\nu}&=~{}~{}\delta\mathcal{T}^{\frac{\omega}{{c_{0}}}}[u_{f}].\end{array}\right.$ (17)
Proposition 4.4.
The solution $u_{f}(\omega,\delta)$ to (15) has the following asymptotic behavior as $\omega,\delta\rightarrow 0$:
$u_{f}(\omega,\delta)=\sum_{j=1}^{N}\frac{a_{j}}{\lvert B_{j}\rvert}1_{B_{j}}+\omega^{2}\sum_{j=1}^{N}\frac{a_{j}}{{c_{r}}^{2}\lvert B_{j}\rvert^{2}}1_{B_{j}}+\delta\sum_{j=1}^{N}a_{j}\left(-\sum_{l=1}^{N}\frac{\operatorname{cap}(\operatorname{{\mathcal{D}}})_{l,j}}{\lvert B_{j}\rvert\lvert B_{l}\rvert^{2}}1_{B_{l}}+\tilde{u}^{j}_{0,1}\right)$
$+\omega\delta\sum_{j=1}^{N}a_{j}\left(\sum_{l=1}^{N}\left(\frac{i}{4\pi\lvert B_{j}\rvert\lvert B_{l}\rvert^{2}{c_{0}}}(\operatorname{cap}(\operatorname{{\mathcal{D}}})J\operatorname{cap}(\operatorname{{\mathcal{D}}}))_{l,j}\right)1_{B_{l}}+\tilde{u}^{j}_{1,1}\right)$
$+\lvert\omega\rvert^{2}\omega i\beta\sum_{j=1}^{N}\frac{\lvert a_{j}\rvert^{2}a_{j}}{\lvert B_{j}\rvert^{4}}1_{B_{j}}+O((\omega^{2}+\delta)^{2}),$
where $J$ denotes the $(N\times N)$-matrix of ones.
Proof.
We make the Ansatz
$u_{f}(\omega,\delta)=\sum_{p,k,l\geq 0}\omega^{p}\delta^{k}\overline{\omega}^{l}u_{p,k,l}.$
Inserting this into (17) and identifying powers of $\delta$ and $\omega$, we obtain the following cascade of equations characterizing the functions $(u_{p,k})_{p,k\geq 0}$:
$-\Delta u_{p,k,l}+\sum_{i=1}^{N}\left(\int_{B_{i}}u_{p,k,l}\operatorname{dvol}\right)1_{B_{i}}$ $=\frac{1}{{c_{r}}^{2}}u_{p-2,k,l}+i\beta\left(\lvert u_{f}\rvert^{2}u_{f}\right)_{p-2,k,l-1}+\sum_{j=1}^{N}a_{j}1_{B_{j}}\delta_{p=0}\delta_{k=0}\delta_{l=0},$
$\frac{\partial u_{p,k,l}}{\partial\nu}$ $=\sum_{n=0}^{p}\frac{1}{{c_{0}}^{n}}\mathcal{T}_{n}[u_{p-n,k-1,l}],$
where $\left(\lvert u_{f}\rvert^{2}u_{f}\right)_{p-2,k,l-1}$ is the $\omega^{p-2}\delta^{k}\overline{\omega}^{l-1}$-component of $\lvert u_{f}\rvert^{2}u_{f}$. For $l=0$, the same computations as in the proof of Proposition 3.8 hold and one obtains that
$u_{p,k,0}=\sum_{j=1}^{N}a_{j}(u_{1,j})_{p,k}.$
Assume $l=1$. Then $u_{0,0,1}=u_{1,0,1}=0$. For $l=1,p=2,k=0,$ one has
$-\Delta u_{2,0,1}+\sum_{i=1}^{N}\left(\int_{B_{i}}u_{2,0,1}\operatorname{dvol}\right)1_{B_{i}}$ $=i\beta\left(\lvert u_{f}\rvert^{2}u_{f}\right)_{0,0,0},$
$\frac{\partial u_{2,0,1}}{\partial\nu}$ $=0.$
Thus,
$\lvert B_{j}\rvert\int_{B_{j}}u_{2,0,1}(x)\operatorname{dvol}=i\beta\int_{B_{j}}\lvert u_{0,0,0}\rvert^{2}u_{0,0,0}\operatorname{dvol},$
and it follows that
$u_{2,0,1}(x)=i\beta\frac{\lvert u_{0,0,0}(x)\rvert^{2}u_{0,0,0}(x)}{\lvert B_{j}\rvert}\quad\text{ for }x\in B_{j}.$
Similarly, it follows that $u_{0,1,1}=u_{1,0,2}=u_{0,0,3}=0$. ∎
This proposition allows to derive the following fundamental theorem of nonlinear subwavelength physics by deriving the asymptotics of the nonlinear subwavelength resonances in terms of $\delta$.
Theorem 4.5.
Let $\{q_{0},v_{2},\ldots,v_{N}\}$ be a basis of eigenvectors of $\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})$, let $V$ be as in Definition 2.3 and denote by $\Pi:\mathbb{C}^{N}\rightarrow\mathbb{C}q_{0}$ the projection along $\operatorname{span}(v_{2},\ldots,v_{N})$ onto $\mathbb{C}q_{0}$. Then the subwavelength resonances $\omega(\delta)=\omega_{0}\sqrt{\delta}+\omega_{1}\delta+O(\delta^{\frac{3}{2}})$ of (6) associated to $q_{0}$ satisfies the following asymptotics
$\omega_{0}\text{ is an eigenvector of }\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})\text{ with eigenvector }q_{0},$
$\omega_{1}=\frac{q_{0}\cdot\Pi\left[\frac{-i}{4\pi{c_{0}}}\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})J\operatorname{cap}(\operatorname{{\mathcal{D}}})q_{0}-\lvert\omega_{0}\rvert^{2}i\beta{c_{r}}^{2}(V^{vol})^{2}\lvert q_{0}\rvert^{2}q_{0}\right]}{2\lVert q_{0}\rVert^{2}},$
where $\lvert q_{0}\rvert^{2}q_{0}$ is to be understood to act component-wise.
Proof.
By Theorem 4.3, the subwavelength resonances of (6) are given by those $\omega(\delta)$ such that there exist $a_{1},\ldots,a_{N}$ such that (15) and (16) are satisfied. Inserting the results of Proposition 4.4 into (16) implies that $\omega(\delta)$ is a subwavelength resonance whenever there exists $q(\delta)\in\mathbb{C}^{N}\setminus\{0\}$ such that
$\begin{split}\left(\omega(\delta)^{2}-\delta\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})+\omega(\delta)\delta\frac{i}{4\pi{c_{0}}}\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})J\operatorname{cap}(\operatorname{{\mathcal{D}}})\right)q(\delta)\quad\quad\quad\quad\quad\quad\\ +\lvert\omega\rvert^{2}\omega i\beta{c_{r}}^{2}(V^{vol})^{2}\lvert q(\delta)\rvert^{2}q(\delta)=0,\end{split}$ (18)
up to an error of order $(\omega(\delta)^{2}+\delta)^{2}$, where $\lvert q(\delta)\rvert^{2}q(\delta)$ is to be understood component-wise. Making the Ansatz $\omega(\delta)=\omega_{0}\sqrt{\delta}+\omega_{1}\delta+O(\delta^{\frac{3}{2}})$ and $q(\delta)=q_{0}+q_{1}\sqrt{\delta}+O(\delta)$ and inserting it into the above equation leads to the following cascade of equations:
$\left(\omega_{0}^{2}-\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})\right)q_{0}=0,$
$\left(\omega_{0}^{2}-\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})\right)q_{1}+\omega_{0}\left(2\omega_{1}q_{0}+\frac{i}{4\pi{c_{0}}}\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})J\operatorname{cap}(\operatorname{{\mathcal{D}}})q_{0}+\lvert\omega_{0}\rvert^{2}i\beta{c_{r}}^{2}(V^{vol})^{2}\lvert q_{0}\rvert^{2}q_{0}\right)=0.$
It then follows that $\omega_{0}$ is an eigenvector of $\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})$ with eigenvector $q_{0}$ and $\omega_{1}$ is given by
$\omega_{1}=\frac{q_{0}\cdot\Pi\left[\frac{-i}{\pi{c_{0}}}\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})J\operatorname{cap}(\operatorname{{\mathcal{D}}})q_{0}-\lvert\omega_{0}\rvert^{2}i\beta{c_{r}}^{2}(V^{vol})^{2}\lvert q_{0}\rvert^{2}q_{0}\right]}{2\lVert q_{0}\rVert^{2}}.$
Remark 4.6.
Note that (18) has some similarities with the discrete Gross-Pitaevskii and Aubry-André-Harper models, which are used as discrete approximations of nonlinear Schrödinger equations (with cubic nonlinearities) in quantum mechanics; see, e.g., [soliton3, soliton4, soliton6]. Such quantum mechanical models are typically based on tight-binding approximations and include only nearest-neighbour interaction terms (long-range interactions are neglected). They correspond to (18) with $\operatorname{cap}^{gen}$ being tridiagonal [flach]. We refer to [weinstein1, weinstein2, weinstein3, localization, pankov] for their analysis.
Remark 4.7.
As in [cochlea], if we consider the scattering of an incident wave $u_{in}$ by a system of nonlinear subwavelength resonators, then based on Theorem 4.5, we can approximate the scattered field $u-u_{in}$ by a few subwavelength modes involving only the nonlinear subwavelength resonances and their associated eigenmodes, so as to retain the physically derived coupling between resonators.
As already mentioned at the beginning of this section, the above procedure is not limited to $g(u)=\lvert\frac{\partial u}{\partial t}\rvert^{2}\frac{\partial u}{\partial t}$ but solely relies on the property that the “Fourier transform” of $g(u)$ is of order $O(\omega^{3})$ as $\omega\rightarrow 0$. The following remark is an illustration of this fact if the order is $O(\omega^{2})$ as $\omega\rightarrow 0$.
Remark 4.8.
If one changes the nonlinearity $i\beta\lvert\omega\rvert^{2}\omega\lvert u\rvert^{2}u$ to $i\beta\omega^{2}\lvert u\rvert^{2}u$, the nonlinearity is of order $\omega^{2}$ thus influencing the leading-order behavior of the subwavelength resonances. To be precise, the cascading system of equations from the proof of Proposition 4.4 is then given by
$-\Delta u_{p,k,l}+\sum_{i=1}^{N}\left(\int_{B_{i}}u_{p,k,l}\operatorname{dvol}\right)1_{B_{i}}$ $=\frac{1}{{c_{r}}^{2}}u_{p-2,k,l}+i\beta\left(\lvert u_{f}\rvert^{2}u_{f}\right)_{p-2,k,l}+\sum_{j=1}^{N}a_{j}1_{B_{j}}\delta_{p=0}\delta_{k=0}\delta_{l=0},$
$\frac{\partial u_{p,k,l}}{\partial\nu}$ $=\sum_{n=0}^{p}\frac{1}{{c_{0}}^{n}}\mathcal{T}_{n}[u_{p-n,k-1,l}],$
and one obtains
$\begin{split}u_{f}(\omega,\delta)=\sum_{j=1}^{N}\frac{a_{j}}{\lvert B_{j}\rvert}1_{B_{j}}+\omega^{2}\left(\sum_{j=1}^{N}\frac{a_{j}}{{c_{r}}^{2}\lvert B_{j}\rvert^{2}}1_{B_{j}}+i\beta\sum_{j=1}^{N}\frac{\lvert a_{j}\rvert^{2}a_{j}}{\lvert B_{j}\rvert^{4}}1_{B_{j}}\right)\quad\quad\quad\quad\quad\quad\\ +\delta\sum_{j=1}^{N}a_{j}\left(-\sum_{l=1}^{N}\frac{\operatorname{cap}(\operatorname{{\mathcal{D}}})_{l,j}}{\lvert B_{j}\rvert\lvert B_{l}\rvert^{2}}1_{B_{l}}+\tilde{u}^{j}_{0,1}\right)\quad\quad\quad\quad\\ +\omega\delta\sum_{j=1}^{N}a_{j}\left(\sum_{l=1}^{N}\left(\frac{-i}{4\pi\lvert B_{j}\rvert\lvert B_{l}\rvert^{2}{c_{0}}}(\operatorname{cap}(\operatorname{{\mathcal{D}}})J\operatorname{cap}(\operatorname{{\mathcal{D}}}))_{l,j}\right)1_{B_{l}}+\tilde{u}^{j}_{1,1}\right)+O((\omega^{2}+\delta)^{2}).\end{split}$
In that case, the subwavelength resonances $\omega(\delta)=\omega_{0}\sqrt{\delta}+O(\delta)$ are characterized by the nonlinear eigenvalue problem
$\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})q_{0}-\omega_{0}^{2}\left(q_{0}+i\beta{c_{r}}^{2}(V^{vol})^{2}\lvert q_{0}\rvert^{2}q_{0}\right)=0,$
where $\lvert q_{0}\rvert^{2}q_{0}$ is to be understood to act component-wise. If $\operatorname{{\mathcal{D}}}$ is connected, then
$\omega_{0}^{2}=\frac{\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})}{1+i\beta{c_{r}}^{2}\frac{\lvert q_{0}\rvert^{2}}{\lvert\operatorname{{\mathcal{D}}}\rvert^{2}}}.$ (19)
5 Numerical experiments: A nonlinear system of two resonators
To illustrate the results from Section 4, some numerical results will be presented in this section. Concretely, we will study the nonlinearity $i\beta\omega^{2}\lvert u\rvert^{2}u$ for the Helmholtz system (6), as discussed in Remark 4.8. The describing equation for the leading-order behavior of the subwavelength modes, is in that case given by
$\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})q_{0}-\omega_{0}^{2}\left(q_{0}+i\beta{c_{r}}^{2}(V^{vol})^{2}\lvert q_{0}\rvert^{2}q_{0}\right)=0,$ (20)
where the entries of the vector $q_{0}$ correspond to the average of the subwavelength resonant solutions and $\lvert q_{0}\rvert^{2}q_{0}$ is to be understood to act component-wise on the entries of $q_{0}$.
In the case where $\operatorname{{\mathcal{D}}}=B_{1}\cup B_{2}$ and $B_{1}$ and $B_{2}$ satisfy certain symmetry conditions, the resonant solutions to the linear system (2) remain resonant solutions of the nonlinear system.
Lemma 5.1.
Assume that $\operatorname{{\mathcal{D}}}$ is a smooth bounded domain with well-separated connected components $B_{1},~{}B_{2}$ with connected boundaries. If $\operatorname{{\mathcal{D}}}=B_{1}\cup B_{2}$ is preserved under the mirror symmetry $P(x_{1},x_{2},x_{3})=(x_{1},x_{2},-x_{3})$ and $P(B_{1})=P(B_{2})$, then, for all $a\in\mathbb{C}$,
$v_{0}=\begin{pmatrix}a\\ a\end{pmatrix}\text{ and }v_{1}=\begin{pmatrix}-a\\ a\end{pmatrix}$
are solutions to equation (20) with
$\omega_{0}=\frac{\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})_{1,1}+\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})_{2,2}}{\left(1+i\beta\frac{{c_{r}}^{2}}{\lvert B_{1}\rvert^{2}}\lvert a\rvert^{2}\right)}\quad\text{ and }\quad\omega_{1}=\frac{\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})_{1,1}-\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})_{2,2}}{\left(1+i\beta\frac{{c_{r}}^{2}}{\lvert B_{1}\rvert^{2}}\lvert a\rvert^{2}\right)},$ (21)
respectively.
Proof.
Since $\operatorname{{\mathcal{D}}}$ satisfies the mirror symmetry condition, it follows that $\operatorname{cap}(\operatorname{{\mathcal{D}}})$ has the form
$\operatorname{cap}(\operatorname{{\mathcal{D}}})=\begin{pmatrix}C_{1}&C_{2}\\ C_{2}&C_{1}\end{pmatrix},$
with $C_{1},C_{2}\in\mathbb{R}$. Furthermore, $V(1,1)=V(2,2)=\lvert B_{1}\rvert^{-1}$. Thus, a basis of eigenvectors of $V\operatorname{cap}(\operatorname{{\mathcal{D}}})$ is given by
$w_{0}=\begin{pmatrix}1\\ 1\end{pmatrix}\text{ and }w_{1}=\begin{pmatrix}-1\\ 1\end{pmatrix},$
with
$\lambda_{0}=\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})_{1,1}+\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})_{2,2}\quad\text{ and }\quad\lambda_{1}=\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})_{1,1}-\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})_{2,2}$
being the respective eigenvalues. It follows that when plugging $v_{0}=aw_{0}$ or $v_{1}=aw_{1}$ into the system (20), one obtains that
$\begin{split}\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})v_{j}-\omega_{0}^{2}\left(v_{j}+i\beta{c_{r}}^{2}(V^{vol})^{2}\lvert a\rvert^{2}v_{j}\right)\\ =\left(\lambda_{j}-\omega_{0}^{2}\left(1+i\beta\frac{{c_{r}}^{2}}{\lvert B_{1}\rvert^{2}}\lvert a\rvert^{2}\right)\right)v_{j}.\end{split}$
It follows that $(v_{j},\omega_{j}(a))$ with
$(\omega_{j}(a))^{2}=\frac{\lambda_{j}}{\left(1+i\beta\frac{{c_{r}}^{2}}{\lvert B_{1}\rvert^{2}}\lvert a\rvert^{2}\right)}$
solves (20). ∎
Remark 5.2.
In the symmetric dimer case, the resonant solutions to the linear system are preserved for all amplitudes of the nonlinearity given in Remark 4.8. Comparing equations (19) and (21), one can observe that the dependence of the subwavelength resonances $\omega(a)$ on the amplitude $a$ is reminiscent of the monomer case discussed in Remark 4.8.
Of course, in the nonlinear setting with a cubic nonlinearity, one expects a third solution. For this solution to occur, the magnitude of the solution modes needs to be sufficiently large. In Figure 2(a), the solutions $q_{0}$ are displayed for a domain $\operatorname{{\mathcal{D}}}$ in the setting of Lemma 5.1. Concretely, the solutions of equation (20) are computed for $B_{1}=B_{0.2}(-\frac{1}{2}e_{3})$ and $B_{2}=B_{0.2}(\frac{1}{2}e_{3})$. The material parameters are chosen as follows ${c_{r}}=1$ and $\beta=0.1$.
Due to the complex nature of the solutions we decided to plot the magnitude of each entry $\lvert(q_{0})_{1}\rvert$, $\lvert(q_{0})_{2}\rvert$ and display the phase of $(q_{0})_{1}/(q_{0})_{2}$ as the respective color of the curve.
One can see the solutions predicted from Lemma 5.1 as the diagonal line in Figure 1(a) (the predicted solutions lie on top of each other). For a sufficiently big magnitude, one sees another solution curve occurring. The amplitudes of the entries $\lvert(q_{0})_{1}\rvert$ and $\lvert(q_{0})_{2}\rvert$ are symmetric and the phase difference is conjugate, as one would expect from a symmetric system. This is made precise in the following lemma.
Lemma 5.3.
Let $\operatorname{{\mathcal{D}}}$ be a smooth bounded domain with well-separated connected components $B_{1},~{}B_{2}$ with connected boundaries. Suppose that $\operatorname{{\mathcal{D}}}=B_{1}\cup B_{2}$ is preserved under the mirror symmetry $P(x_{1},x_{2},x_{3})=(x_{1},x_{2},-x_{3})$ and $P(B_{1})=P(B_{2})$. Then, if $(\omega,q_{0})$ is a solution of (20), then
$\left(\omega,p_{0}\right)\quad\text{ with }\quad p_{0}=\begin{pmatrix}(q_{0})_{2}\\ (q_{0})_{1}\end{pmatrix}$
is also a solution of (20) and the phase of $(p_{0})_{1}/(p_{0})_{2}$ is conjugate to the phase of $(q_{0})_{1}/(q_{0})_{2}$.
Proof.
When $\operatorname{{\mathcal{D}}}$ satisfies the above symmetry, it follows that $\operatorname{cap}^{gen}(\operatorname{{\mathcal{D}}})$ is symmetric. Plugging $(\omega,p_{0})$ into equation (20), one observes that the first component of (20) is given by the second component when plugging in $(\omega,q_{0})$ and the second is equal to the first. Thus, both are equal to zero and $(\omega,p_{0})$ is a solution to (20). ∎
In Figure 2(b) and Figure 2(c), the solutions for asymmetric systems are computed. The domain $B_{2}$ is changed to $B_{0.21}(\frac{1}{2}e_{3})$ and $B_{0.22}(\frac{1}{2}e_{3})$, respectively. For $\lvert(q_{0})_{1}\rvert$, $\lvert(q_{0})_{2}\rvert$ small enough, one observes that the solution curves are close to the solutions of the linear setting, as one would expect from linearization techniques. Due to the asymmetry of the system, the solutions $q_{0}$ of the linear system have entries of different magnitudes $\lvert(q_{0})_{1}\rvert\not=\lvert(q_{0})_{2}\rvert$, which makes the two subspaces visible (light blue and red curves no longer lie on top of each other, as in Figure 1(a)). For sufficiently large magnitudes, a third solution appears. In the symmetric setting, the third solution did not interfere with the ‘linear’ solutions predicted by Lemma 5.1, when the symmetry is broken, the third solution interferes with the first two. This can be seen in the splitting of the light blue curve.
Figure 1: In these plots the solution curves of equation (20) for different choices of $\operatorname{{\mathcal{D}}}$ are plotted. From left to right, $\operatorname{{\mathcal{D}}}=B_{0.2}\left(-\frac{1}{2}e_{3}\right)\cup B_{0.2}\left(\frac{1}{2}e_{3}\right)$, $\operatorname{{\mathcal{D}}}=B_{0.2}\left(-\frac{1}{2}e_{3}\right)\cup B_{0.21}\left(\frac{1}{2}e_{3}\right)$ and $\operatorname{{\mathcal{D}}}=B_{0.2}\left(-\frac{1}{2}e_{3}\right)\cup B_{0.22}\left(\frac{1}{2}e_{3}\right)$, respectively. The remaining parameters are given by $\beta=0.1$ and ${c_{r}}=1$. The curves trace the magnitude of the entries of different solutions $q_{0}$ to equation (20). The colors account for the complex valued solutions $q_{0}\in\mathbb{C}^{2}$ and depict the phase of $(q_{0})_{1}/(q_{0})_{2}$.
The respective subwavelength resonances of the nonlinear setting are displayed in Figure 2. The subwavelength resonances of the symmetric setting associated to Figure 1(a) are depicted in Figure 2(a). In the main plot the value of subwavelength resonances are plotted in the complex plane. In the inlet plot the subwavelength solution with the colors of the different solution curves corresponding to the different curves of the subwavelength resonances are plotted. In Figure 2(a), one observes that the curve of the subwavelength resonances which are associated to the third solution curve, touches the curve of the subwavelength resonances associated to the solution $q_{0}=(a,a)^{\top},a\in\mathbb{C}$, and the superscript $\top$ denotes the transpose. It is precisely that curve that is mixing with the nonlinear solutions for sufficiently large amplitudes in the asymmetric settings, as can be seen from Figures 1(b) and 2(b), and Figures 1(c) and 2(c), respectively.
Figure 2: In this figure the subwavelength resonances associated to the settings discussed in the caption of Figure 1 are displayed. The resonances are displayed as curves in the complex plane. The color of each curve is associated to the color of a solution curve depicted in the inlet plots.
6 Conclusion and outlook
In this paper, we have provided for the first time a discrete approximation to the nonlinear resonance problem for finite systems of subwavelength resonators. Our discrete approximation allows not only deriving the effect of weak nonlinearities on the linear resonances but more importantly characterizes the soliton-like solutions which are induced by stronger nonlinearities. It would be very interesting, firstly, to rigorously derive modal decompositions for scattered solutions by nonlinear systems of resonators and, secondly, to analyze the stability of the soliton-like resonances and their associated eigenmodes. Our results in this paper can be extended to periodic systems of resonators. We plan in forthcoming works to elucidate the interplay between nonlinearities and subwavelength resonances, in particular the nonlinearity-induced topological transitions in such systems.
Acknowledgments
This work was supported in part by the Swiss National Science Foundation grant number 200021–200307. T.K. would like to thank Tianwei Yu and Wouter Tonnon for the numerous discussions.
[…]